# Revealing Human Internal Attention Patterns from Gameplay Analysis for Reinforcement Learning

### v2

@@ 4.1 Visual Analysis of Attention Maps
Figure 5: Binarized attention map comparison: Example scenes (unseen states) with binarized TIOA (upper) and CTR attention map (lower) overlay. CTR human attention maps more closely resemble the TIOA than the agent attention. They likely represent human interal attention patterns as they include parts of the TIOA map, objects nearby, and highly task-relevant objects (see annotations).
@@
In Fig. 5, a scene is presented for each game with overlays of the TIOA and CTR maps. The selected scenes are intended to represent typical characteristics of the TIOA and CTR maps. Animations are provided in the supplemental materials. To adjust for differences between human and agent attention differences identified in Sec. 3 we base the chosen activation rates $r_{\Psi}$ off the 95% of action prediction accuracy threshold from 4, with a lower limit of $0.04$ and an upper limit of $0.16$ resulting in the values shown in Tab. 1. In Fig. 5 for better visualization, we use a stricter upper limit of $0.08$.
@@
   Enduro    0.16    0.097
   Freeway    0.096    0.04
   MsPacman    0.16    0.077
   Seaquest    0.143    0.04
   SpaceInvaders    0.16    0.04
   Riverraid    0.16    0.041
@@
It will be qualitatively analyzed for each game whether the TIOA and attention maps differ, and if the CTR attention maps plausibly resemble internal attention. Again, the TIOA maps serve as a point of comparison as no true label of the internal attention is available. Generally, the TIOA maps are expected to overlap better with the human CTR attention maps but not fully.
@@
In Freeway, the agent consistently attends to most cars and the player’s avatar (a chicken crossing the road). In contrast, human CTR attention alternates between two distinct modes: when walking, the focus is mostly on the avatar (seen in 6), and when walking, attention shifts toward the road ahead and cars (seen in 5). The 2 modes can also be seen in quantiative analysis in the following Sec. 4.2. The pattern aligns with human gaze behavior, where attention is primarily directed at the path ahead and objects near it. The human generally focuses stronger on the line the avatar is walking than the agent. This is predicted by the CTR network and can be seen clearly in the heatmap as well.
For MsPacman and Seaquest, the TIOA predicts human gaze primarily around the avatar character and a few opponents or collectibles. The human CTR attention map likely reflects a pattern of interal attention, especially in MsPacman, where the similarities are strongest. In the shown scene, the CTR attention map focuses on a relevant nearby ghost which is also looked at by the human avatar, but rather at the edge of focus area. The agent’s CTR map, on the other hand, extends attention to opponents and smaller items farther away. The agent’s attention is stronger along the horizontal line along the avatar’s vertical position. Notably, heatmaps show that both the human CTR and TIOA maps attend to various positions in the game environment.
@@ 4.2 Similarity Analysis and Histograms of Attention Maps
We observe that both human and agent CTR attention maps outperform the baseline of $1.0$. For all games, the human CTR maps reach a higher alignment score averaging at $3.26$ compared to the agent CTR maps, which reach $2.49$. The alignment scores vary across games, with the highest human-agent gap in Freeway and MsPacman, indicating that human attention patterns in these games are more closely aligned with TIOA predictions. Space Invaders and Riverraid show the smallest differences. For Freeway, we observe a long tail end of the histogram, this reflects the mode where the human is predicted to focus on the road ahead as described in Sec. 3. Interestingly, the alignment score per bin also spikes here as well, underlying that this pattern exhibits consistency with the TIOA predictions. Overall, these results demonstrate that human CTR attention maps consistently exhibit a strong similarity to the TIOA maps across all games, indicating that the CTR network can capture human internal attention patterns.
5 Evaluation III: Human Internal Attention-Guided Reinforcement Learning
@@
Figure 9: Comparison of learning curves of the Dueling DDQN agent with PER for the plain agent, the human internal attention-guided agent (CTR), and the inverted attention agent. Three runs are averaged per agent type, after performing a rolling average of 200 steps per run. Shaded area indicates standard deviation. Overall, CTR-masked agents slightly outperform plain ones, showing stable learning.
@@
Achieving human scores in Atari games is a challenging benchmark for RL agents, especially under limited training steps. Even algorithms such as Rainbow Deep Q-Learning (DQN) [32] that combine several improvements over normal DQN fail to reach mean human-normalized performance within 10 million environment steps, while human-outperforming methods like Agent57 [33] are highly tailored for Atari. Guo et al. [23] found that RL agents (here, using Proximal Policy Optimization, PPO), when they fail, are more likely to neglect regions attended to by humans. We investigate whether the human internal attention maps can improve learning performance, by making RL agents focus on objects that are task-relevant for humans. This represents an alternative approach to imitation learning, as it aims to transfer not just actions but the underlying attentional strategies derived from those actions.
@@
As a strong baseline, we use DQN with the modifications of double DQN (DDQN) [34], dueling DQN [35], and proportional prioritized experience replay (PER) [36]. Dueling DQN improves learning stability by separating value and advantage streams, double DQN mitigates overestimation bias by using a separate target network, and PER allows the agent to learn more effectively from rare but significant events. These enhancements are implemented using a custom-modified version of the stable baselines 3 (SB3) [37] DQN implementation. The policy network receives the feature representation $\mathcal{F}_{\mathrm{p}}$ as input. We evaluate three agent variants: (1) a plain agent using features from a pre-trained, frozen autoencoder ($\mathcal{F}_{\mathrm{p}}=\mathcal{F}$), (2) a human internal attention-guided agent, where features are masked by the attention map $\bm{\Psi}$ to focus on task-relevant objects ($\mathcal{F}_{\mathrm{p}}=\bm{\Psi}\odot\mathcal{F}$), and (3) an inverted attention agent, where features are masked by the inverted CTR map ($\mathcal{F}_{\mathrm{p}}=(1-\bm{\Psi})\odot\mathcal{F}$). The comparison to the inverted map is done similarly to [38], providing a control baseline to evaluate the usefulness of the attention maps, especially since selective attention does not always guarantee performance gains [39]. For each agent (1-3), we perform three runs of different seeds. For all experiments, we use $16\%$ average activation rate in the CTR attention maps (meaning $84\%$ average activation for the inverted maps).
@@
Fig. 9 shows the learning curves for the three agent variants, denoting the reward obtained during the actual training episodes. The human internal attention-guided agent outperforms the plain agent for Enduro, Freeway and MsPacman, achieves similar performance for SpaceInvaders and slightly worse for Seaquest and Riverraid. The inverted attention agent performs worse in most games, especially in Seaquest where almost no learning progress is made. Unintuitively, its performance in Freeway is similar to the plain agent. The CTR-masked agents also appear to learn quite stably, especially in Enduro and MsPacman.
Figure 10: Comparison of episodic scores of the Dueling DDQN agent with PER for the plain agent, the human internal attention-guided agent (CTR), and the inverted attention agent, including the random agent score. Three runs are averaged per agent type. Shaded area indicates standard deviation. Overall, CTR-masked agents consistently achieve higher episodic scores compared to plain agents.
Fig. 10 shows the average episodic scores for the three agent variants after training. Here, the agents are evaluated under an exploration rate of $\epsilon=0.01$ for 150 episodes (50 per agent of different seeds), meaning they exploit their learned policies while still choosing a random action in $1\%$ of the cases. An overall average relative score is determined by
@@
is determined using random scores [^5] Random agent scores: Enduro: 0.0, Freeway: 0.0, MsPacman: 307.3, Seaquest: 68.4, SpaceInvaders: 148.0, Riverraid: 1338.5, taken from [33] and plain scores as the baselines. The CTR-masked agent achieves highest episodic scores for MsPacman, Riverraid, and SpaceInvaders. It clearly outperforms the inverted attention agent in all games except Freeway, where all three agent types perform similarly. For the average relative score, the CTR-masked agents achieve $1.10$ whereas the inverted attention agents underperform with a score of $0.68$. The human internal attention-guided agents therefore reach $10\%$ higher scores than the baseline Dueling DDQN agent with PER. This shows that the derived attention maps can slightly improve RL agents performance, limiting the policy network’s available feature space to the $16\%$ predicted to be most relevant for humans.
@@ 6 Conclusions
Our study presents a novel method to derive human internal attention patterns from gameplay data alone, leveraging offline attention methods from reinforcement learning (RL). By introducing the contextualized, task-relevant (CTR) attention network, we generate attention maps from human and agent gameplay, compared against a temporally-integrated overt attention (TIOA) model based on eye-tracking data. After qualitative and quantitative evaluation, these maps are applied for attention-guided reinforcement learning.
Evaluation across six Atari games demonstrates that human CTR attention maps exhibit higher alignment towrad the TIOA maps compared to those generated for agents. Overall, the human CTR maps capture key regions of the TIOA maps, as well as adjacent objects and highly task-relevant features, providing evidence that the CTR network reveals internal human attention patterns. Applied to RL, CTR-masked agents achieved 10% higher scores than baseline using only 16% of the feature space, demonstrating improved learning efficiency and performance, offering an alternative to imitation learning approaches.
@@
Our results represent a first step toward predicting human internal attention from interactions alone, without relying on additional data such as brain activity. This approach can deepen understanding of human-agent attention differences and contribute to human-inspired autonomous agents. Future work could explore more advanced strategies for human internal attention-guided learning, aiding the development of explainable, interpretable, and robust autonomous systems.
@@
Supplementary information
We provide supplementary animations of the CTR and TIOA attention maps, namely the binarized maps shown in Fig. 5, the smoothed maps shown in Fig. 6, as well as the maps binarized at average activation of 16% as used in the RL experiments in Sec. 5.
@@
The authors did not receive support from any organization for the submitted work.

### v3

@@ 4.1 Visual Analysis of Attention Maps
Figure 6: Binarized attention map comparison: Example scenes (unseen states) with binarized TIOA (upper) and CTR attention map (lower) overlay. CTR human attention maps more closely resemble the TIOA than the agent attention. They likely represent human internal attention patterns as they include parts of the TIOA map, objects nearby, and highly decision-relevant objects (see annotations).
@@
In Fig. 6, a scene is presented for each game with overlays of the TIOA and CTR maps. The selected scenes are intended to represent typical characteristics of the TIOA and CTR maps. Animations are provided in the supplemental materials. To adjust for differences between human and agent attention patterns identified in Sec. 3 we base the chosen activation rates $r_{\Psi}$ off the 97% of action prediction accuracy threshold from 4, with a lower limit of $0.04$ and an upper limit of $0.16$ resulting in the values shown in Tab. 1. In Fig. 6 for better visualization, we use a stricter upper limit of $0.08$.
@@
   Enduro    0.065    0.078
   Freeway    0.04    0.04
   MsPacman    0.112    0.066
   Seaquest    0.054    0.04
   SpaceInvaders    0.059    0.04
   Riverraid    0.064    0.048
@@
It will be qualitatively analyzed for each game whether the TIOA and attention maps differ, and if the CTR attention maps plausibly resemble internal attention patterns. Again, the TIOA maps serve as a point of comparison as no true label for the decision-relevant features exists. Generally, the TIOA maps are expected to overlap better with the human CTR attention maps but not fully.
@@
In Freeway, the agent consistently attends to most cars and the player’s avatar (a chicken crossing the road). In contrast, human CTR attention alternates between two distinct modes: when walking, the focus is mostly on the avatar (seen in 7), and when walking, attention shifts toward the road ahead and cars (seen in 6). The 2 modes can also be seen in quantiative analysis in the following Sec. 4.2. The pattern aligns with human gaze behavior, where attention is primarily directed at the path ahead and objects near it. The human generally focuses stronger on the line the avatar is walking than the agent. This is predicted by the CTR network and can be seen clearly in the heatmap as well.
For MsPacman and Seaquest, the TIOA predicts human gaze primarily around the avatar character and a few opponents or collectibles. The human CTR attention map likely reflects a pattern of internal attention, especially in MsPacman, where the similarities are strongest. In the shown scene, the CTR attention map focuses on a relevant nearby ghost which is also looked at by the human avatar, but rather at the edge of focus area. The agent’s CTR map, on the other hand, extends attention to opponents and smaller items farther away. The agent’s attention is stronger along the horizontal line along the avatar’s vertical position. Notably, heatmaps show that both the human CTR and TIOA maps attend to various positions in the game environment.
@@ 4.2 Similarity Analysis and Histograms of Attention Maps
We observe that both human and agent CTR attention maps outperform the baseline of $1.0$. Across all games, the human CTR maps achieve a higher average alignment score of $3.26$, while the agent CTR maps reach an average of $2.49$. To assess the statistical significance of this difference, we performed a one-sided t-test using the average alignment score for each game as an independent sample. The result indicates a statistically significant difference in favor of the human CTR maps, with a $p$-value of $0.00031$. The alignment scores vary across games, with the highest human-agent gap in Freeway and MsPacman, indicating that human attention patterns in these games are more closely aligned with TIOA predictions. Space Invaders and Riverraid show the smallest differences. For Freeway, we observe a long tail end of the histogram, this reflects the mode where the human is predicted to focus on the road ahead as described in Sec. 3. Interestingly, the alignment score per bin also spikes here as well, underlying that this pattern exhibits consistency with the TIOA predictions. Overall, these results demonstrate that human CTR attention maps consistently exhibit a strong similarity to the TIOA maps across all games, indicating that the CTR network can reveal human internal attention patterns. While the CTR network extracts generalized patterns rather than exact momentary attention states, the combination of high action prediction accuracy under masking and statistically significant gaze alignment provides empirical support that these patterns reflect the features causally relevant to human decision-making in these games.
5 Evaluation III: Human Attention-Guided Reinforcement Learning
@@
Figure 10: Comparison of learning curves of the Dueling DDQN agent with PER for the plain agent, the human attention-guided agent (CTR), the TIOA-guided agent, and the inverted attention agent. Three runs are averaged per agent type, after performing a rolling average of 200 steps per run. Shaded area indicates standard deviation. Overall, CTR-masked agents slightly outperform plain ones, showing stable learning, while TIOA-masked agents perform poorly.
@@
Achieving human-level scores in Atari games is a challenging benchmark for RL agents, especially under limited training steps. Even algorithms such as Rainbow Deep Q-Learning (DQN) [hessel2018rainbow] that combine several improvements over normal DQN fail to reach mean human-normalized performance within 10 million environment steps, while human-outperforming methods like Agent57 [badia2020agent57] are highly tailored for Atari. Guo et al. [guo2021machine] found that RL agents (here, using Proximal Policy Optimization, PPO), when they fail, are more likely to neglect regions attended to by humans, and that agent attention becomes more similar to human gaze when reducing the discount factor from 0.99 to 0.9, suggesting that human attention patterns can guide RL agents toward more human-like behavior. We investigate whether the human attention patterns can improve learning performance, by making RL agents focus on objects that are decision-relevant for humans. This represents an alternative approach to imitation learning, as it aims to transfer not only actions but the underlying attentional strategies derived from those actions.
@@
As a strong baseline, we use DQN with the modifications of double DQN (DDQN) [van2016deep], dueling DQN [pmlr-v48-wangf16], and proportional prioritized experience replay (PER) [schaul2015prioritized]. Dueling DQN improves learning stability by separating value and advantage streams, double DQN mitigates overestimation bias by using a separate target network, and PER allows the agent to learn more effectively from rare but significant events. These enhancements are implemented using a custom-modified version of the stable baselines 3 (SB3) [JMLR:v22:20-1364] DQN implementation. The policy network receives the feature representation $\mathcal{F}_{\mathrm{p}}$ as input. We evaluate four agent variants: (1) a plain agent using features from a pre-trained, frozen autoencoder ($\mathcal{F}_{\mathrm{p}}=\mathcal{F}$), (2) a human attention-guided agent, where features are masked by the CTR attention map $\bm{\Psi}$ to focus on decision-relevant objects ($\mathcal{F}_{\mathrm{p}}=\bm{\Psi}\odot\mathcal{F}$), with attention values below 0.1 set to zero to ensure no information leakage at low attention values, (3) a TIOA-guided agent, where features are masked by binarized TIOA gaze prediction maps ($\mathcal{F}_{\mathrm{p}}=\hat{\bm{\Gamma}}_{\mathrm{bin}}\odot\mathcal{F}$), and (4) an inverted attention agent, where features are masked by the inverted CTR map ($\mathcal{F}_{\mathrm{p}}=(1-\bm{\Psi})\odot\mathcal{F}$). The comparison to the inverted map is done similarly to [itaya2021visual], providing a control baseline to evaluate the usefulness of the attention maps, especially since selective attention does not always guarantee performance gains [yuezhang2018initial]. The TIOA-guided agent allows us to assess whether CTR attention provides additional value beyond overt attention alone for feature-level masking. For each agent (1-4), we perform three runs of different seeds. For all experiments, we use $16\%$ average activation rate in the CTR attention maps (meaning $84\%$ average activation for the inverted maps). For the TIOA-guided agents, the binarization threshold is set to achieve the same $16\%$ average activation rate, enabling a fair comparison of feature selection ability at matched sparsity levels. We apply hard masking for both CTR and TIOA to directly compare their effectiveness at selecting relevant features, though other approaches to incorporate gaze information (e.g., as auxiliary input channels or through soft masking) are possible and may perform differently.
@@
Fig. 10 shows the learning curves for the four agent variants, denoting the reward obtained during the actual training episodes. The human CTR-guided agent outperforms the plain agent for Enduro, Freeway and MsPacman, achieves similar performance for SpaceInvaders and slightly worse for Seaquest and Riverraid. The inverted attention agent performs worse in most games, especially in Seaquest where almost no learning progress is made. Unintuitively, its performance in Freeway is similar to the plain agent. The CTR-masked agents also appear to learn quite stably, especially in Enduro and MsPacman. The TIOA-masked agents perform poorly across all games, worse than the inverted CTR agents except for Seaquest.
Figure 11: Comparison of episodic scores of the Dueling DDQN agent with PER for the plain agent, the human CTR-masked agent, the TIOA-masked agent, and the inverted attention agent, including the random agent score. Three runs are averaged per agent type. CTR-masked agents in average achieve higher episodic scores, while TIOA-masked agents perform poorly.
Fig. 11 shows the average episodic scores for the four agent variants after training. Here, the agents are evaluated under an exploration rate of $\epsilon=0.01$ for 150 episodes (50 per seed) per agent type, meaning they exploit their learned policies while still choosing a random action in $1\%$ of the cases. An overall average relative score is determined by
@@
is determined using random scores [^5] Random agent scores: Enduro: 0.0, Freeway: 0.0, MsPacman: 307.3, Seaquest: 68.4, SpaceInvaders: 148.0, Riverraid: 1338.5, taken from [badia2020agent57] and plain scores as the baselines. The CTR-masked agent achieves highest episodic scores for MsPacman, Riverraid, and SpaceInvaders. It clearly outperforms the inverted attention agent in all games except Freeway, where the plain, CTR-masked, and inverted attention agents perform similarly, while TIOA-masked agents perform worse. The TIOA-masked agents perform poorly across all games, even outperformed by inverterted CTR-masked agent except in Seaquest.
For the average relative score, the CTR-masked agents achieve $1.10$, while the inverted attention agents underperform with a score of $0.68$, and the TIOA-masked agents achieve $0.52$. Thus, the human CTR-guided agents reach $10\%$ higher scores than the baseline Dueling DDQN agent with PER. To assess statistical significance, we conducted one-sided t-tests using each game’s average relative score as an independent sample per model. The CTR-masked agents significantly outperformed both the TIOA-masked agents ($p=0.0024$) and the inverted attention agents ($p=0.0162$). While the comparison with the plain, baseline agents did not reach statistical significance ($p=0.238$), the CTR-masked agents still achieved 10% higher scores on average across the tested games. Overall, although the performance improvement over baseline is not statistically significant across all games, these findings indicate a positive trend suggesting that CTR attention masks may improve RL agent performance.
The lower performance of TIOA-masked agents suggests that binarizing gaze maps does not effectively select decision-relevant feature channels at the encoded level. In contrast, CTR attention learns to select feature channels that capture decision-relevant patterns, demonstrating its value beyond overt attention. We note that this comparison evaluates hard feature selection at matched sparsity; other approaches to incorporate gaze information may perform differently.
@@ 6 Conclusions
Our study presents a novel method to derive human attention patterns from gameplay data alone, leveraging offline attention methods from reinforcement learning (RL). By introducing the contextualized, task-relevant (CTR) attention network, we generate attention maps from human and agent gameplay, compared against a temporally-integrated overt attention (TIOA) model based on eye-tracking data. After qualitative and quantitative evaluation, these maps are applied for attention-guided reinforcement learning.
Evaluation across six Atari games demonstrates that human CTR attention maps identify features relevant for action prediction, and reach statistically significantly higher alignment toward the TIOA maps compared to those generated for agents. Overall, the human CTR maps capture key regions of the TIOA maps, as well as adjacent objects and highly decision-relevant features, providing evidence that the CTR network reveals human internal attention patterns. Applied to RL, CTR-masked agents achieved 10% higher scores than baseline over the tested games, using only 16% of the feature space. This hints at improved learning efficiency and performance and offers an alternative to imitation learning approaches. They further statistically significantly outperform TIOA-masked agents and inverted CTR-masked agents, reaching a $2.1$ and $1.6$ times higher episodic score respectively.
@@
Our results represent a first step toward predicting human internal attention patterns from interactions alone, without relying on additional data such as brain activity. This approach can deepen understanding of human-agent attention differences and contribute to human-inspired autonomous agents. Future work could explore more advanced strategies for human attention-guided learning, aiding the development of explainable, interpretable, and robust autonomous systems.
@@
Supplementary information We provide supplementary animations of the CTR and TIOA attention maps, namely the binarized maps shown in Fig. 6, the smoothed maps shown in Fig. 7, as well as the maps binarized at average activation of 16% as used in the RL experiments in Sec. 5.
@@
Funding: The authors did not receive support from any organization for the submitted work.
Competing Interests: The authors declare no competing interests.